OpenStudy (anonymous):

which chemical formula is correct? Ag9CH3CO2)2, K2(CN), Fe(OH)2, or MgHCO3

jhonyy9 (jhonyy9):

probably MgHCO3

OpenStudy (anonymous):

no fe(oh)2 is the correct one mg has valency of 2 its mg(hco3)2

OpenStudy (anonymous):

using oxidation numbers: Mg = 2+ H = +1 C = 4+ 0 = 3 x -2 = -6 total for HCO3 = -1 so we need 2 * HCO3 to balance with the Mg

OpenStudy (anonymous):

ash plz close these questions when you are happy by clicking the good question box